OpenStudy (anonymous):

12x^2-1=-x solve the equationa dn check your solution

OpenStudy (anonymous):

first put it in the standard form ax^2+bx+c =0 and post it here

OpenStudy (anonymous):

??

OpenStudy (anonymous):

add x to both sides of the equation and post what you get.

OpenStudy (anonymous):

i dont get what you mean so yeahhh

OpenStudy (anonymous):

have you done algebra before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol then how come you have to do these questions?

OpenStudy (anonymous):

oh. okay. Then you really shouldn't be doing this sort of problem. Let me post a couple of links so that you can first learn some basic algebra and then come back here.

OpenStudy (anonymous):

this is my first year doing algebra and i have a 27% and have to retake it next yuear

OpenStudy (anonymous):

oh. You really should learn how to do these questions. They're not hard ^^ Here's your answer with explanations: 12x² - 1 = -x [[move everything to one side and get the equation into the general form ax²+bx+c =0]] 12x² + x - 1 = 0 [[You're looking for 2 numbers, m and n, with the conditions that m+n = b and mn = ac. So in this case, two numbers that add to be 1 and multiply to be -12. 4 and -3 work, so split x into 4x and -3x]] 12x² + 4x - 3x - 1 = 0 [[Factorize by grouping in pairs]] 4x(3x + 1) - (3x + 1) = 0 [[Factorise (3x+1) out]] (4x - 1)(3x + 1) = 0 [[Solve]] x = 1/4 or x = -1/3